OpenStudy (anonymous):

-2[square root of the 16(16)(-2) - (-2)^3(-1)] over (2 - 6)^2

OpenStudy (anonymous):

@Venny

OpenStudy (anonymous):

Is this what you're asking?

OpenStudy (anonymous):

-(Squareroot of 123/2)/4 would be the exact answer

OpenStudy (anonymous):

no

OpenStudy (anonymous):

@hhunny13

OpenStudy (anonymous):

@math-geek

OpenStudy (anonymous):

see it below...one moment,,

OpenStudy (anonymous):

\[\frac{ -2\sqrt{16*16-2-8*-1} }{ (-4)^2 }=\frac{ -2\sqrt{256-2+8} }{ 16 }=\frac{ -\sqrt{262} }{ 8 }\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@zaibali.qasmi

OpenStudy (anonymous):

yes,..???

OpenStudy (anonymous):

its not 16 times 16

OpenStudy (anonymous):

this is the quation

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

holding...

OpenStudy (anonymous):